A ball is tossed straight upwards with an initial
Answer #1

3)

initial velocity=v=-30m/s( upwards)

initial height=20 cm

velocity =v+gt=-30+gt

distance =S=20+vt+0.5gt^2=20-30t+4.9t^2

(b)velocity after hitting ground=v=sqrt(900+2*9.8*20)=35.9 m.s

(c) velocity after bounce=8.98

height=v^2/2g=4.11 m